Express the function g, defined by g(x)=1/√(x+√x) as a composition of three functions.
1
Expert's answer
2021-09-16T06:16:01-0400

Solution;

One possible way ;

If you let;

"f(x)=\sqrt{x}"

"g(x)=\sqrt{x^2+x}"

"h(x)=\frac1x"

First,obtain g of f(x);

"g o f=g(f(x))=\sqrt{(f(x))^2+f(x)}=\sqrt{(\sqrt{x})^2+\sqrt{x}}=\sqrt{x+\sqrt{x}}"

Now we plug g(f(x)) into h(x);

"F(x)=h (gof)=h(g(f(x)))"

"F(x)=\frac{1}{g(f(x))}" ="\frac{1}{\sqrt{x+\sqrt{x}}}"

Hence the expression of the function is three functions could be expressed as;

"f(x)=\sqrt{x}" ,"g(x)=\sqrt{x^2+x}" , "h(x)=\frac1x" ,"F(x)=h" o "g" o "f"